OpenStudy (anonymous):

Solve: 4- (x+3) = 8- 5(2x-3)

OpenStudy (anonymous):

whr r u stuck?

OpenStudy (anonymous):

i got 9x=16 and then 1 and 7/9 i just wasnt sure if it was right or not

OpenStudy (anonymous):

show me how

OpenStudy (precal):

use the draw feature to show your steps so we can help you figure out where you made a mistake

OpenStudy (anonymous):

k 4- (x+3) = 8- 5(2x-3) 7-x=8-10x+15 7-x=23-10x 9x=16

OpenStudy (anonymous):

4- (x+3) = 8- 5(2x-3) 1-x=8-10x+15 ; not 7

OpenStudy (anonymous):

omg thanks! would it be 9x=22 then and x= 2 and 4/9

OpenStudy (anonymous):

awsumm, correct